OpenStudy (anonymous):

\[\int_{0}^{\pi/2} cos^{2}\theta d\theta\]

OpenStudy (anonymous):

use this formula \[\cos^2 \theta=\frac{1+\cos 2 \theta}{2}\]

OpenStudy (anonymous):

why can't I just say \[\frac 1 3 cos^3 \theta\]? or is that illegal?

OpenStudy (anonymous):

that is illegal because \[\frac{d}{d \theta} (\frac{1}{3} \cos^3 \theta)= -\sin \theta \ \cos^2 \theta\]

OpenStudy (anonymous):

true...thanks

OpenStudy (anonymous):

yw